OpenStudy (fswiney2):

Anna and Belle are asked to solve –2x – 15 = 6x + 9. Identify where one of them made an error. Anna Belle –2x – 15 = 6x + 9 –2x – 15 –6x = 6x + 9 –6x –8x – 15 = 9 –8x – 15 + 15 = 9 + 15 –8x = 24 x = –3 –2x – 15 = 6x + 9 –2x – 15 + 2x = 6x + 9 + 2x –15 = 8x + 9 –15 + 9 = 8x + 9 + 9 –6 = 8x

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Hint: Look at these steps –15 = 8x + 9 –15 + 9 = 8x + 9 + 9 <---how does 9 + 9 = 0 –6 = 8x